OpenStudy (crashonce):

ABCD is a quadrilateral. E, F are the midpoints of AC and BD respectively. Prove that AB^2 + BC^2 + CD^2 +DA^2=AC^2+BD^2+4EF^2

OpenStudy (crashonce):

@SolomonZelman @jim

OpenStudy (crashonce):

@jim_thompson5910 my bad

OpenStudy (crashonce):

@zzr0ck3r

OpenStudy (zzr0ck3r):

never taken a Geometry class.

OpenStudy (crashonce):

oh by the way, do all this using only algebra

OpenStudy (zzr0ck3r):

and fact about geometry I am sure....

OpenStudy (crashonce):

@iambatman @iGreen @TuringTest

OpenStudy (crashonce):

@mathmate

OpenStudy (crashonce):

@wio

OpenStudy (crashonce):

@Compassionate

OpenStudy (crashonce):

@surjithayer you can only use algebra

OpenStudy (mathmate):

Start with drawing a diagram of the quadrilateral ABCD, with the mid points of AC and BD named E and F respectively. Without loss of generality (WLOG), we can place A at (0,0), and B on the x-axis. We end up with the following coordinates: \(A(0,0)\) \(B(b,0)\) \(C(c,r)\) \(D(d,s)\) \(E(\frac{c}{2}, \frac{r}{2})\) \(F(\frac{d+b}{2}, \frac{s}{2})\)

OpenStudy (mathmate):



OpenStudy (mathmate):

I will do the first part, using distance between two points: \(AB^2+BC^2+CD^2+DA^2\) \(=(b^2)+((c-b)^2+r^2)+((d-c)^2+(s-r)^2)+(d^2+s^2)\) \(=2s^2-2rs+2r^2+2d^2-2cd+2c^2-2bc+2b^2\) You can complete the same exercise for the right-hand side: \(AC^2+BD^2+4EF^2~=~...\) and show that the two parts are equal.

OpenStudy (mathmate):

E is the mid-point of AC, so E((0+c)/2, (0+r)/2)=(c/2,r/2) F is the mid-point of BD, so F((b+d)/2, (0+s)/2)=((b+d)/2, s/2)

OpenStudy (mathmate):

THe algebra might be a little messy when it comes to the length of \(EF^2\), but it is rewarding at the end of the calculations.

OpenStudy (crashonce):

Yep got it thanks so much

OpenStudy (mathmate):

You're welcome! :)